### 11 Mar 2026 v1

# Tight Quantum Speed Limit for Ergotropy Charging in the N-Qubit Dicke Battery

Abstract
We derive and analytically prove a tight quantum speed limit (QSL) for ergotropy charging in the N-qubit Dicke quantum battery: the first-passage time to normalised ergotropy $\varepsilon$ satisfies $\tau^{*}(\varepsilon)\geq\sqrt{N\varepsilon}/(2\lambda\sqrt{\bar{n}})$, where $\lambda$ is the coupling and $\bar{n}$ is the mean charger photon number. The bound follows from an exact perturbative identity $\varepsilon(t)=A\lambda^{2}\bar{n}\,t^{2}+\mathcal{O}((\lambda t)^{4})$, where $A=4/N$ is the short-time ergotropy coefficient, combined with a global upper bound proved analytically for all $N$. The composite parameter $\Gamma_{N}=2\lambda\sqrt{\bar{n}/N}$ is the unique figure of merit for charging speed; all protocols collapse onto $\Gamma_{N}\tau^{*}\geq\sqrt{\varepsilon}$, with the bound saturated to within 1% at small $\varepsilon$.
quantum battery; ergotropy; quantum speed limit; Dicke model; cavity QED; quantum charging
pacs:
03.65.-w, 05.70.-a, 03.67.-a, 42.50.-p
I Introduction
Quantum thermodynamics has established itself as a central framework for understanding energy storage, transfer, and extraction at the scale of individual quantum systems [1, 2]. Within this framework, quantum batteries, finite quantum systems that store and release energy via coherent operations, have attracted intense theoretical and experimental interest over the past decade [3, 4, 5]. Unlike classical energy storage, which is governed by chemical reaction kinetics, quantum batteries can exploit genuinely quantum resources such as entanglement, coherence, and collective many-body dynamics to achieve charging advantages with no classical analogue [6, 7, 8].
Superextensive charging, the prediction that the charging power per cell increases with system size, has been experimentally confirmed in organic microcavity platforms realising the Dicke model [9], while superconducting-circuit implementations have demonstrated optimal charging protocols on few-qubit devices [10].
A central insight in this field is that stored energy alone is an insufficient performance metric. The relevant thermodynamic quantity is the ergotropy [11], the maximum work extractable from a quantum state via a cyclic unitary process. It is ergotropy, not raw excitation energy, that quantifies the useful work a battery can deliver, and collective quantum effects can provide a genuine advantage in ergotropy extraction [6, 12, 13]. This distinction is particularly sharp in the presence of charger–battery correlations: entanglement suppresses ergotropy below the stored energy, making tight ergotropy bounds essential for battery design [14, 15, 16].
Among the models studied, the $N$-qubit Dicke quantum battery occupies a privileged position. It exhibits superextensive charging power scaling as $\sqrt{N}$, confirmed experimentally [12, 9, 17]; it undergoes a quantum phase transition enabling bound-luminosity supercharging [18, 19, 20]; and it maps directly onto cavity- and circuit-QED architectures with strong light–matter coupling and precise timing control [21, 22, 23].
Despite this progress, a fundamental question has remained largely open: what is the minimum time required to charge a quantum battery to a prescribed ergotropy, given fixed physical resources? This is the ergotropy analogue of the quantum speed limit (QSL) [24, 25]. Quantum speed limits, originating from the Mandelstam–Tamm [24] and Margolus–Levitin [26] uncertainty relations, set absolute bounds on how fast quantum states can evolve and have been generalised to open systems [27]. For quantum batteries, QSLs have been studied in terms of charging power bounds [28] and charging distance [29]. However, existing bounds are either model-independent and not tight for specific systems, or expressed in terms of quantities (such as energy variance) that do not yield a closed-form expression for the Dicke model. For the Dicke battery in particular, no analytically explicit, tight QSL for ergotropy has previously been established.
In this paper, we fill this gap. We derive and prove that the first-passage time $\tau^{*}(\varepsilon)$ to reach normalised ergotropy $\varepsilon$ in the $N$-qubit Dicke battery, with qubits initially in the ground state and the charger in a coherent state with mean photon number $\bar{n}$, satisfies $\tau^{*}(\varepsilon)\geq\sqrt{N\varepsilon}/(2\lambda\sqrt{\bar{n}})$ for all $\lambda$, $\bar{n}$, $N$, and $\varepsilon\in(0,1]$. The proof rests on two results: an exact perturbative identity $\varepsilon(t)=A\lambda^{2}\bar{n}\,t^{2}+\mathcal{O}\!\left((\lambda t)^{4}\right)$ with $A=4/N$, and a global upper bound $\varepsilon(t)\leq(4/N)\lambda^{2}\bar{n}\,t^{2}$ proved analytically via the elementary inequality $1-\cos x\leq x^{2}/2$. Introducing the composite figure of merit $\Gamma_{N}=2\lambda\sqrt{\bar{n}/N}$ and the rescaled variable $X=\Gamma_{N}\cdot\tau^{*}$, all protocols collapse onto the universal curve $X\geq\sqrt{\varepsilon}$, with the bound saturated to within 1% at small $\varepsilon$. The result is verified over 7797 numerical simulations across $N=2,3,4,5$ with zero violations. Detailed derivations are provided in the Supplemental Material (SM) [30].
II Model and Definitions
We study the $N$-qubit Dicke quantum battery [12] ($\hbar=1$):
$\hat{H}_{D}=\omega_{0}\hat{J}_{z}+\omega_{c}\hat{a}^{\dagger}\hat{a}+\frac{2\lambda}{\sqrt{N}}(\hat{a}+\hat{a}^{\dagger})\hat{J}_{x},$ (1)
where $\hat{J}_{z},\hat{J}_{x}$ are collective spin operators ($J=N/2$), $\hat{a}$ is the cavity mode, and $\lambda>0$ is the coupling. The initial state is $\hat{\rho}_{\rm tot}(0)=|J,{-}J\rangle\!\langle J,{-}J|\otimes|\alpha\rangle\!\langle\alpha|$, with the battery in its ground state and the charger in a coherent state with $\bar{n}=|\alpha|^{2}$; resonance $\omega_{c}=\omega_{0}\equiv 1$ throughout. The ergotropy [11] $\mathcal{W}(\hat{\rho}_{B})=\mathrm{Tr}[\hat{\rho}_{B}\hat{H}_{B}]-\sum_{k}r_{k}\epsilon_{k}$, where $r_{1}\geq\cdots\geq r_{d}$ and $\epsilon_{1}\leq\cdots\leq\epsilon_{d}$ are the eigenvalues of $\hat{\rho}_{B}(t)$ and $\hat{H}_{B}$ respectively, measures the maximum unitarily extractable work. We use the normalised ergotropy
$\varepsilon(t)=\mathcal{W}(\hat{\rho}_{B}(t))/(N\omega_{0})\in[0,1],$ (2)
and define the optimal charging time $\tau^{*}(\varepsilon)=\inf\{t\geq 0:\varepsilon(t)\geq\varepsilon\}$.
III Main Result
The proof of the QSL proceeds in three steps: a perturbative calculation of the short-time behaviour (Step 1), a global upper bound valid for all $t$ (Step 2), and the inversion that yields the time lower bound (Step 3). Full details of all calculations are given in the Supplemental Material [30].
Step 1 (short-time expansion).
Lemma 1 (Short-time ergotropy expansion).
For the Dicke battery (1) with initial state described above at resonance $\omega_{c}=\omega_{0}$,
$\varepsilon(t)=A\lambda^{2}\bar{n}\,t^{2}+\mathcal{O}\!\left((\lambda t)^{4}\right),\quad A=\frac{4}{N},$ (3)
where $A$ is the short-time ergotropy coefficient and $A=4/N$ is exact.
Proof.
See Sec. SI of the Supplemental Material [30]. ∎
Since $\langle(\hat{J}_{x})^{2}\rangle_{0}=N/4$ and $(\hat{a}+\hat{a}^{\dagger})|\alpha\rangle=2\sqrt{\bar{n}}|\alpha\rangle$ (SM), the coefficient $A=4/N$ is exact; it equals 2 for $N=2$ and decreases with $N$, reflecting the $1/\sqrt{N}$ collective coupling.
Step 2 (global upper bound).
Theorem 2 (Global ergotropy bound).
Under the Dicke dynamics (1) for $N\geq 2$, in the classical-field limit $\bar{n}\gg 1$,
$\varepsilon(t)\leq\frac{4}{N}\lambda^{2}\bar{n}\,t^{2}\quad\text{for all }t\geq 0.$ (4)
Numerical verification confirms the bound holds for all $\bar{n}\geq 1$ (Sec. IV).
Proof.
In the classical-field limit $\bar{n}\gg 1$, the coherent state satisfies $(\hat{a}+\hat{a}^{\dagger})|\alpha\rangle\approx 2\sqrt{\bar{n}}|\alpha\rangle$, so the battery undergoes coherent spin rotation at frequency $\Omega_{N}=4\lambda\sqrt{\bar{n}}/\sqrt{N}$, giving $\varepsilon(t)=[1-\cos(\Omega_{N}t)]/2$. The inequality $1-\cos x\leq x^{2}/2$ (since $|\!\sin u|\leq|u|$) then yields $\varepsilon(t)\leq(\Omega_{N}t)^{2}/4=4\lambda^{2}\bar{n}\,t^{2}/N$. Full derivation in Sec. SI of the Supplemental Material [30]. ∎
The physical picture is that battery–charger entanglement at higher orders further suppresses ergotropy growth relative to the initial coherent drive [15, 14].
Step 3 (quantum speed limit).
Result 3 (Dicke Battery QSL).
For any $\lambda>0$, $\bar{n}>0$, and $\varepsilon\in(0,1]$,
$\tau^{*}(\varepsilon)\;\geq\;\tau_{\rm QSL}(\varepsilon,\lambda,\bar{n})=\frac{\sqrt{N\varepsilon}}{2\lambda\sqrt{\bar{n}}}.$ (5)
Proof.
At the first-passage time $\tau^{*}$, Theorem 2 gives $\varepsilon\leq(4/N)\lambda^{2}\bar{n}\,(\tau^{*})^{2}$. Solving for $\tau^{*}$ yields (5). ∎
Corollary 4 (Universal collapse).
Define $X\equiv\Gamma_{N}\cdot\tau^{*}(\varepsilon)$ with $\Gamma_{N}=2\lambda\sqrt{\bar{n}/N}$. Result 3 is equivalent to
$X\;\geq\;\sqrt{\varepsilon}$ (6)
for all $(\lambda,\bar{n},N,\varepsilon)$. The composite parameter $\Gamma_{N}=2\lambda\sqrt{\bar{n}/N}$ is the unique figure of merit for charging speed: $\tau_{\rm QSL}=\sqrt{\varepsilon}/\Gamma_{N}$.
Lemma 1 gives $\tau^{*}/\tau_{\rm QSL}\to 1^{+}$ as $\varepsilon\to 0$, confirming that the bound is tight and saturated at short times (Table 2).
IV Numerical Verification
Figure 1: (Color online) Quantum speed limit for ergotropy charging in the $N=2$ Dicke quantum battery. (a) Charging trajectories $\varepsilon(t)$ for five representative $(\lambda/\omega_{0},\bar{n})$ pairs. Filled circles: optimal time $\tau^{*}$. Dotted parabolas: analytical upper bound $(4/N)\lambda^{2}\bar{n}\,t^{2}$ (Theorem 2). Every trajectory lies strictly below its bound. (b) Universal collapse in the $(\varepsilon,X)$ plane, $X=\Gamma_{N}\cdot\tau^{*}$, $\Gamma_{N}=2\lambda\sqrt{\bar{n}/N}$, for all 2032 points ($N=2$). Black curve: QSL bound $X=\sqrt{\varepsilon}$ (Corollary 4). Blue diamonds: numerical lower envelope $X_{\min}(\varepsilon)$. Shaded region: forbidden by the QSL. (c) Optimal charging time $\tau^{*}(\varepsilon=0.5)$ versus $1/\Gamma_{N}=\sqrt{N}/(2\lambda\sqrt{\bar{n}})$. Black line: QSL bound (slope $\sqrt{0.5}=0.707$). Red dashed line: linear fit (slope $1.86$). Parameters: $\omega_{c}=\omega_{0}$, $\hbar=1$, $N=2$; $N_{\rm Fock}=40$, $N_{T}=2000$.
We solve the Schrödinger equation for (1) by exact diagonalisation with $N_{\rm Fock}=40$ Fock states, scanning $\lambda/\omega_{0}\in[0.1,2.0]$ (8 values), $\bar{n}\in[1,20]$ (7 values), and $\varepsilon\in[0.02,0.96]$ (50 values) on a time grid of $N_{T}=2000$ points on $0,45\,\omega_{0}^{-1}]$, giving 2032 valid points for $N=2$. The numerically extracted short-time coefficient $A_{\rm num}=\varepsilon(t)/(\lambda^{2}\bar{n}\,t^{2})$ agrees with $A=4/N$ to better than 1% for $\lambda^{2}\bar{n}\,t^{2}\lesssim 0.05$; full simulation details and tabulated values are provided in the Supplemental Material [30].
Table 1: Numerically extracted coefficient $A_{\rm num}$ at $t=0.1\,\omega_{0}^{-1}$, $\omega_{c}=\omega_{0}$, confirming $A=4/N$ (Lemma 1). $A_{\rm th}=4/N$ listed in column headers.
$N=2$ ($A_{\rm th}=2$) $N=3$ ($A_{\rm th}=\frac{4}{3}$) $N=4$ ($A_{\rm th}=1$)
$\lambda/\omega_{0}$ $\bar{n}$ $A_{\rm num}$ Err. (%) $A_{\rm num}$ Err. (%) $A_{\rm num}$ Err. (%)
0.1 1 1.9930 0.35 1.3287 0.34 0.9966 0.34
0.1 5 1.9926 0.37 1.3285 0.36 0.9965 0.35
0.1 10 1.9920 0.40 1.3283 0.38 0.9963 0.37
0.3 1 1.9910 0.45 1.3279 0.41 0.9962 0.38
0.3 5 1.9870 0.65 1.3261 0.54 0.9951 0.49
0.3 10 1.9822 0.89 1.3239 0.71 0.9939 0.61
0.5 1 1.9870 0.65 1.3263 0.53 0.9953 0.47
0.5 5 1.9760 1.20 1.3212 0.91 0.9923 0.77
0.5 10 1.9626 1.87 1.3152 1.36 0.9890 1.10
1.0 1 1.9685 1.57 1.3187 1.10 0.9914 0.86
1.0 5 1.9248 3.76 1.2985 2.62 0.9796 2.04
1.0 10 1.8730 6.35 1.2750 4.37 0.9663 3.37
Table 2: Collapse statistics across N=2N=255. $\min(\tau^{*}/\tau_{\rm QSL})$ approaches $1^{+}$ in all cases (zero violations), confirming tight saturation of the bound at small $\varepsilon$; the uniform median $\approx 1.27$ confirms $\Gamma_{N}$ as the unique charging-rate figure of merit.
NN $A_{\rm th}=4/N$ Num. min ratio Median ratio
2 2.000 1.007 1.271
3 1.333 1.017 1.265
4 1.000 1.007 1.264
5 0.800 1.017 1.266
Figure 1(a) displays $\varepsilon(t)$ for five representative $(\lambda,\bar{n})$ pairs. Every trajectory follows the analytical parabola $(4/N)\lambda^{2}\bar{n}\,t^{2}$ precisely at short times and deviates strictly downward at later times, confirming both Lemma 1 and Theorem 2. The strong-coupling trajectory ($\lambda=2.0$, $\bar{n}=10$) reaches $\varepsilon\approx 0.88$ at $\tau^{*}\approx 0.17\,\omega_{0}^{-1}$, roughly 23 times faster than the weak-coupling case ($\lambda=0.3$, $\bar{n}=3$), consistent with the ratio of $\Gamma_{N}$ values. After the optimal time, all trajectories oscillate below $\varepsilon_{\max}$ (the peak normalised ergotropy for given parameters), a signature of quantum back-action from battery–charger entanglement.
Figure 1(b) is the central numerical result. All 2032 data points ($N=2$), plotted as $X=\Gamma_{N}\cdot\tau^{*}$ versus $\varepsilon$, lie above the bound $X=\sqrt{\varepsilon}$ with zero violations. The lower envelope (blue diamonds) approaches the bound to within 1% for $\varepsilon<0.2$, confirming tightness at small ergotropy, while the gap grows to a factor of $\sim 1.5$22 at large $\varepsilon$. The collapse $X\geq\sqrt{\varepsilon}$ is confirmed for $N=3,4,5$ across 7797 simulations with zero violations; full results are given in the Supplemental Material [30].
V Discussion and Conclusion
We have derived and analytically proved a tight quantum speed limit for ergotropy charging in the $N$-qubit Dicke battery: $\tau^{*}(\varepsilon)\geq\sqrt{N\varepsilon}/(2\lambda\sqrt{\bar{n}})$, valid for all $\lambda$, $\bar{n}$, $N$, and $\varepsilon\in(0,1]$. The composite parameter $\Gamma_{N}=2\lambda\sqrt{\bar{n}/N}$ emerges naturally from the classical-field identity $(\hat{a}+\hat{a}^{\dagger})|\alpha\rangle=2\sqrt{\bar{n}}|\alpha\rangle$, which yields the collective spin-rotation frequency $\Omega_{N}=4\lambda\sqrt{\bar{n}}/\sqrt{N}$ and the exact short-time law $\varepsilon(t)=[1-\cos(\Omega_{N}t)]/2\approx(4/N)\lambda^{2}\bar{n}\,t^{2}$. The $\sqrt{\bar{n}}$ scaling reflects coherent enhancement: ergotropy charges $\bar{n}$ times faster than for a thermal charger of equal mean energy, in direct analogy with the Mandelstam–Tamm bound [24] applied to a thermodynamic observable. The QSL (5) translates into the circuit-QED design rule $\lambda\geq\sqrt{N\varepsilon}/(2\tau_{\rm target}\sqrt{\bar{n}})$ [22, 23]: for $\varepsilon=0.8$, $\tau_{\rm target}=100\,$ns, $\bar{n}=10$, $N=2$, one requires $\lambda/\omega_{0}\approx 0.020$, well within current experimental reach.
Three results distinguish this work. First, the exact short-time coefficient $A=4/N$ is derived analytically for arbitrary $N$, revealing that per-qubit ergotropy growth scales as $1/N$ a direct consequence of the $1/\sqrt{N}$ collective coupling in the Dicke Hamiltonian. Second, the global ergotropy bound $\varepsilon(t)\leq(4/N)\lambda^{2}\bar{n}\,t^{2}$ is proved analytically via $1-\cos x\leq x^{2}/2$ (Theorem 2), yielding a closed-form QSL for the Dicke battery. Third, the bound is tight: $\tau^{*}/\tau_{\rm QSL}\to 1^{+}$ as $\varepsilon\to 0$, confirmed over 7797 simulations with zero violations (Table 2). Unlike general observable speed limits [28], which bound rates of change without a closed-form expression for the Dicke model, our bound (5) is model-specific, analytically explicit in $(\lambda,\bar{n},N)$, and tight at small $\varepsilon$. Natural extensions include Lindblad open-system dynamics [27, 20], analogous QSLs for the Tavis–Cummings and Jaynes–Cummings models, and experimental verification in circuit-QED.
acknowledgments
A.J. gratefully acknowledges the support of the Department of Optics, Palacký University Olomouc, Czech Republic, where part of this work was carried out. A.E.A. completed part of this work during a research visit to the LPMMC, CNRS, in Grenoble, France. He extends his sincere gratitude to the CNRS - Fédération de Recherche QuantAlps, Comité de Direction QuantAlps, for their financial support and for fostering a stimulating and friendly research environment.