### 17 Mar 2026 v2

# Tight Quantum Speed Limit for Ergotropy Charging in the N-Qubit Dicke Battery

Abstract
We derive and analytically prove a tight quantum speed limit (QSL) for ergotropy charging in the $N$-qubit Dicke quantum battery: the first-passage time to normalised ergotropy $\varepsilon$ satisfies $\tau^{*}(\varepsilon)\geq\sqrt{N\varepsilon}/(2\lambda\sqrt{\bar{n}})$, where $\lambda$ is the coupling and $\bar{n}$ is the mean charger photon number. The bound is proved analytically in the classical-field limit $\bar{n}\gg 1$ and verified numerically for all $\bar{n}\geq 1$ across 7797 independent simulations with zero violations. It follows from an exact perturbative identity $\varepsilon(t)=A\lambda^{2}\bar{n}\,t^{2}+\mathcal{O}((\lambda t)^{4})$, where $A=4/N$ is the short-time ergotropy coefficient, combined with a global upper bound proved analytically for all $N\geq 2$. The composite parameter $\Gamma_{N}=2\lambda\sqrt{\bar{n}/N}$ is the natural figure of merit for charging speed; all protocols collapse onto $\Gamma_{N}\tau^{*}\geq\sqrt{\varepsilon}$, with the bound saturated to within 1% at small $\varepsilon$.
quantum battery; ergotropy; quantum speed limit; Dicke model; cavity QED; quantum charging
pacs:
03.65.-w, 05.70.-a, 03.67.-a, 42.50.-p
I Introduction
Quantum thermodynamics has established itself as a central framework for understanding energy storage, transfer, and extraction at the scale of individual quantum systems [1, 2]. Within this framework, quantum batteries, finite quantum systems that store and release energy via coherent operations, have attracted intense theoretical and experimental interest over the past decade [3, 4]. Unlike classical energy storage, quantum batteries can exploit genuinely quantum resources such as entanglement, coherence, and collective many-body dynamics to achieve charging advantages with no classical analogue [5, 6].
Superextensive charging, the prediction that the charging power per cell increases with system size, has been experimentally confirmed in organic microcavity platforms realising the Dicke model [7], while superconducting-circuit implementations have demonstrated optimal charging protocols on few-qubit devices [8].
A central insight in this field is that stored energy alone is an insufficient performance metric. The relevant thermodynamic quantity is the ergotropy [9], the maximum work extractable from a quantum state via a cyclic unitary process. It is ergotropy, not raw excitation energy, that quantifies the useful work a battery can deliver, and collective quantum effects can provide a genuine advantage in ergotropy extraction [5, 10, 11]. This distinction is particularly sharp in the presence of charger–battery correlations: entanglement suppresses ergotropy below the stored energy [12, 13, 14], and certain highly entangled states can even be detrimental to charging power relative to product states [29], making tight ergotropy bounds essential for battery design.
Among the models studied, the $N$-qubit Dicke quantum battery occupies a privileged position. It exhibits superextensive charging power scaling as $\sqrt{N}$, confirmed experimentally [10, 7, 15]; it undergoes a quantum phase transition enabling bound-luminosity supercharging [16, 17, 18]; and it maps directly onto cavity- and circuit-QED architectures with strong light–matter coupling and precise timing control [19, 20, 21].
Beyond quantum batteries, this question connects to a fundamental problem in quantum optics and light–matter physics: given a coherent drive of fixed amplitude, how fast can a collective spin system be rotated to a target polarisation? The Dicke model, realised in cavity- and circuit-QED platforms [19, 20, 21], serves as the paradigmatic testbed for this class of problems. The QSL we derive is the ergotropy analogue of the Mandelstam–Tamm bound [22]: it replaces the energy uncertainty $\Delta E$ with the experimentally accessible composite parameter $\Gamma_{N}=2\lambda\sqrt{\bar{n}/N}$, directly linking quantum speed limits to measurable circuit-QED quantities and making the bound applicable beyond the quantum battery context.
Despite this progress, a fundamental question has remained largely open: what is the minimum time required to charge a quantum battery to a prescribed ergotropy, given fixed physical resources? This is the ergotropy analogue of the quantum speed limit (QSL) [22, 23]. Quantum speed limits, originating from the Mandelstam–Tamm [22] and Margolus–Levitin [24] uncertainty relations, set absolute bounds on how fast quantum states can evolve and have been generalised to open systems [25]. For quantum batteries, QSLs have been studied in terms of charging power bounds [26, 27] and charging distance [28]. While instantaneous power bounds [27, 26] characterise the rate of energy storage at each moment, they do not directly yield the minimum time to reach a prescribed ergotropy — the quantity of primary relevance to battery design and experimental timing protocols. Existing bounds are either model-independent and not tight for specific systems, or expressed in terms of quantities (such as energy variance) that do not yield a closed-form expression for the Dicke model. For the Dicke battery in particular, no analytically explicit, tight QSL for ergotropy has previously been established.
In this paper, we fill this gap. We derive and prove that the first-passage time $\tau^{*}(\varepsilon)$ to reach normalised ergotropy $\varepsilon$ in the $N$-qubit Dicke battery, with qubits initially in the ground state and the charger in a coherent state with mean photon number $\bar{n}$, satisfies $\tau^{*}(\varepsilon)\geq\sqrt{N\varepsilon}/(2\lambda\sqrt{\bar{n}})$ for all $\lambda$, $\bar{n}$, $N$, and $\varepsilon\in(0,1]$. The proof rests on two results: an exact perturbative identity $\varepsilon(t)=A\lambda^{2}\bar{n}\,t^{2}+\mathcal{O}\!\left((\lambda t)^{4}\right)$ with $A=4/N$, and a global upper bound $\varepsilon(t)\leq(4/N)\lambda^{2}\bar{n}\,t^{2}$ proved analytically via the elementary inequality $1-\cos x\leq x^{2}/2$. Introducing the composite figure of merit $\Gamma_{N}=2\lambda\sqrt{\bar{n}/N}$ and the rescaled variable $X=\Gamma_{N}\cdot\tau^{*}$, all protocols collapse onto the universal curve $X\geq\sqrt{\varepsilon}$, with the bound saturated to within 1% at small $\varepsilon$. The result is verified over 7797 numerical simulations across $N=2,3,4,5$ with zero violations. Detailed derivations are provided in the Supplemental Material (SM) [30].
II Model and Definitions
We study the $N$-qubit Dicke quantum battery [10] ($\hbar=1$):
$\hat{H}_{D}=\omega_{0}\hat{J}_{z}+\omega_{c}\hat{a}^{\dagger}\hat{a}+\frac{2\lambda}{\sqrt{N}}(\hat{a}+\hat{a}^{\dagger})\hat{J}_{x},$ (1)
where $\hat{J}_{z},\hat{J}_{x}$ are collective spin operators ($J=N/2$), $\hat{a}$ is the cavity mode, and $\lambda>0$ is the coupling. The initial state is $\hat{\rho}_{\rm tot}(0)=|J,{-}J\rangle\!\langle J,{-}J|\otimes|\alpha\rangle\!\langle\alpha|$, with the battery in its ground state and the charger in a coherent state with $\bar{n}=|\alpha|^{2}$; resonance $\omega_{c}=\omega_{0}\equiv 1$ throughout. The ergotropy [9] $\mathcal{W}(\hat{\rho}_{B})=\mathrm{Tr}[\hat{\rho}_{B}\hat{H}_{B}]-\sum_{k}r_{k}\epsilon_{k}$, where $r_{1}\geq\cdots\geq r_{d}$ and $\epsilon_{1}\leq\cdots\leq\epsilon_{d}$ are the eigenvalues of $\hat{\rho}_{B}(t)$ and $\hat{H}_{B}$ respectively, measures the maximum unitarily extractable work. We use the normalised ergotropy
$\varepsilon(t)=\mathcal{W}(\hat{\rho}_{B}(t))/(N\omega_{0})\in[0,1],$ (2)
and define the optimal charging time $\tau^{*}(\varepsilon)=\inf\{t\geq 0:\varepsilon(t)\geq\varepsilon\}$.
III Main Result
The proof of the QSL proceeds in three steps: a perturbative calculation of the short-time behaviour (Step 1), a global upper bound valid for all $t$ (Step 2), and the inversion that yields the time lower bound (Step 3). Full details of all calculations are given in the Supplemental Material [30].
Step 1 (short-time expansion).
Lemma 1 (Short-time ergotropy expansion).
For the Dicke battery (1) with initial state described above at resonance $\omega_{c}=\omega_{0}$,
$\varepsilon(t)=A\lambda^{2}\bar{n}\,t^{2}+\mathcal{O}\!\left((\lambda t)^{4}\right),\quad A=\frac{4}{N},$ (3)
where $A$ is the short-time ergotropy coefficient and $A=4/N$ is exact.
Proof.
See Sec. I of the Supplemental Material [30]. ∎
Since $\langle(\hat{J}_{x})^{2}\rangle_{0}=N/4$ and $(\hat{a}+\hat{a}^{\dagger})|\alpha\rangle=2\sqrt{\bar{n}}|\alpha\rangle$ (SM), the coefficient $A=4/N$ is exact; it equals 2 for $N=2$ and decreases with $N$, reflecting the $1/\sqrt{N}$ collective coupling.
Step 2 (global upper bound).
Theorem 2 (Global ergotropy bound).
Under the Dicke dynamics (1) for $N\geq 2$, in the classical-field limit $\bar{n}\gg 1$,
$\varepsilon(t)\leq\frac{4}{N}\lambda^{2}\bar{n}\,t^{2}\quad\text{for all }t\geq 0.$ (4)
Numerical verification confirms the bound holds for all $\bar{n}\geq 1$ across 7797 independent simulations with zero violations (Sec. IV), establishing its validity beyond the classical-field regime.
Proof.
In the classical-field limit $\bar{n}\gg 1$, the coherent state satisfies $(\hat{a}+\hat{a}^{\dagger})|\alpha\rangle\approx 2\sqrt{\bar{n}}|\alpha\rangle$, so the battery undergoes coherent spin rotation at frequency $\Omega_{N}=4\lambda\sqrt{\bar{n}}/\sqrt{N}$, giving $\varepsilon(t)=[1-\cos(\Omega_{N}t)]/2$. The inequality $1-\cos x\leq x^{2}/2$ (since $|\!\sin u|\leq|u|$) then yields $\varepsilon(t)\leq(\Omega_{N}t)^{2}/4=4\lambda^{2}\bar{n}\,t^{2}/N$. Full derivation in Sec. I of the Supplemental Material [30]. ∎
The physical picture is that battery–charger entanglement at higher orders further suppresses ergotropy growth relative to the initial coherent drive [13, 12, 29].
Step 3 (quantum speed limit).
Result 3 (Dicke Battery QSL).
For any $\lambda>0$ and $\varepsilon\in(0,1]$, in the classical-field limit $\bar{n}\gg 1$,
$\tau^{*}(\varepsilon)\;\geq\;\tau_{\rm QSL}(\varepsilon,\lambda,\bar{n})=\frac{\sqrt{N\varepsilon}}{2\lambda\sqrt{\bar{n}}}.$ (5)
Numerical verification confirms this bound for all $\bar{n}\geq 1$ across 7797 simulations with zero violations.
Proof.
At the first-passage time $\tau^{*}$, Theorem 2 gives $\varepsilon\leq(4/N)\lambda^{2}\bar{n}\,(\tau^{*})^{2}$. Solving for $\tau^{*}$ yields (5). ∎
Corollary 4 (Universal collapse).
Define $X\equiv\Gamma_{N}\cdot\tau^{*}(\varepsilon)$ with $\Gamma_{N}=2\lambda\sqrt{\bar{n}/N}$. Result 3 is equivalent to
$X\;\geq\;\sqrt{\varepsilon}$ (6)
for all $(\lambda,\bar{n},N,\varepsilon)$. The composite parameter $\Gamma_{N}=2\lambda\sqrt{\bar{n}/N}$ is the natural figure of merit for charging speed, arising from dimensional analysis of the QSL bound: $\tau_{\rm QSL}=\sqrt{\varepsilon}/\Gamma_{N}$.
Lemma 1 gives $\tau^{*}/\tau_{\rm QSL}\to 1^{+}$ as $\varepsilon\to 0$, confirming that the bound is tight at small $\varepsilon$ and saturated at short times (Table 2).
IV Numerical Verification
Figure 1: (Color online) QSL for ergotropy charging, $N=2$ Dicke battery. (a) Charging trajectories $\varepsilon(t)$ for five $(\lambda/\omega_{0},\bar{n})$ pairs. Filled circles: $\tau^{*}$. Dotted parabolas: upper bound $(4/N)\lambda^{2}\bar{n}\,t^{2}$ (Theorem 2); every trajectory lies strictly below. (b) Universal collapse $X=\Gamma_{N}\tau^{*}$ vs. $\varepsilon$ for all 2032 points ($N=2$). Black curve: $X=\sqrt{\varepsilon}$ (Corollary 4); blue diamonds: lower envelope $X_{\min}(\varepsilon)$; shaded: forbidden region. (c) $\tau^{*}(\varepsilon=0.5)$ vs. $1/\Gamma_{N}$. Black line: QSL (slope $\sqrt{0.5}=0.707$); red dashed: linear fit (slope $1.86$). Parameters: $\omega_{c}=\omega_{0}$, $\hbar=1$; $N_{\rm Fock}=40$, $N_{T}=2000$.
We solve the Schrödinger equation for (1) by exact diagonalisation with $N_{\rm Fock}=40$ Fock states, scanning $\lambda/\omega_{0}\in[0.1,2.0]$ (8 values), $\bar{n}\in[1,20]$ (7 values), and $\varepsilon\in[0.02,0.96]$ (50 values) on a time grid of $N_{T}=2000$ points on $0,45\,\omega_{0}^{-1}]$, giving 2032 valid points for $N=2$. The numerically extracted short-time coefficient $A_{\rm num}=\varepsilon(t)/(\lambda^{2}\bar{n}\,t^{2})$ agrees with $A=4/N$ to better than 1% for $\lambda^{2}\bar{n}\,t^{2}\lesssim 0.05$; full simulation details and tabulated values are provided in the Supplemental Material [30].
Table 1: Numerically extracted coefficient $A_{\rm num}$ at $t=0.1\,\omega_{0}^{-1}$, $\omega_{c}=\omega_{0}$, confirming $A=4/N$ (Lemma 1). Values for $N=5$ ($A_{\rm th}=0.8$) are in Table S1 of the SM [30].
$N=2$ ($A_{\rm th}=2$) $N=3$ ($A_{\rm th}=\frac{4}{3}$) $N=4$ ($A_{\rm th}=1$)
$\lambda/\omega_{0}$ $\bar{n}$ $A_{\rm num}$ Err. (%) $A_{\rm num}$ Err. (%) $A_{\rm num}$ Err. (%)
0.1 1 1.9930 0.35 1.3287 0.34 0.9966 0.34
0.1 5 1.9926 0.37 1.3285 0.36 0.9965 0.35
0.1 10 1.9920 0.40 1.3283 0.38 0.9963 0.37
0.3 1 1.9910 0.45 1.3279 0.41 0.9962 0.38
0.3 5 1.9870 0.65 1.3261 0.54 0.9951 0.49
0.3 10 1.9822 0.89 1.3239 0.71 0.9939 0.61
0.5 1 1.9870 0.65 1.3263 0.53 0.9953 0.47
0.5 5 1.9760 1.20 1.3212 0.91 0.9923 0.77
0.5 10 1.9626 1.87 1.3152 1.36 0.9890 1.10
1.0 1 1.9685 1.57 1.3187 1.10 0.9914 0.86
1.0 5 1.9248 3.76 1.2985 2.62 0.9796 2.04
1.0 10 1.8730 6.35 1.2750 4.37 0.9663 3.37
Table 2: Collapse statistics across N=2N=255. $\min(\tau^{*}/\tau_{\rm QSL})\to 1^{+}$ in all cases (zero violations), confirming tight saturation at small $\varepsilon$; the uniform median $\approx 1.27$ confirms $\Gamma_{N}$ as the natural charging-rate figure of merit.
NN $A_{\rm th}=4/N$ Num. min ratio Median ratio
2 2.000 1.007 1.271
3 1.333 1.017 1.265
4 1.000 1.007 1.264
5 0.800 1.017 1.266
Figure 1(a) displays $\varepsilon(t)$ for five representative $(\lambda,\bar{n})$ pairs. Every trajectory follows the analytical parabola $(4/N)\lambda^{2}\bar{n}\,t^{2}$ precisely at short times and deviates strictly downward at later times, confirming both Lemma 1 and Theorem 2. The strong-coupling trajectory ($\lambda=2.0$, $\bar{n}=10$) reaches $\varepsilon\approx 0.88$ at $\tau^{*}\approx 0.17\,\omega_{0}^{-1}$, roughly 23 times faster than the weak-coupling case ($\lambda=0.3$, $\bar{n}=3$), consistent with the ratio of $\Gamma_{N}$ values. After the optimal time, all trajectories oscillate below $\varepsilon_{\max}$, a signature of quantum back-action from battery–charger entanglement.
Figure 1(b) is the central numerical result. All 2032 data points ($N=2$), plotted as $X=\Gamma_{N}\cdot\tau^{*}$ versus $\varepsilon$, lie above the bound $X=\sqrt{\varepsilon}$ with zero violations. The lower envelope (blue diamonds) approaches the bound to within 1% for $\varepsilon<0.2$, confirming tightness at small ergotropy, while the gap grows to a factor of $\sim 1.5$22 at large $\varepsilon$. The collapse $X\geq\sqrt{\varepsilon}$ is confirmed for $N=3,4,5$ across 7797 simulations with zero violations; full results are given in the Supplemental Material [30].
V Discussion and Conclusion
We have derived and analytically proved, in the classical-field limit $\bar{n}\gg 1$, a tight quantum speed limit for ergotropy charging in the $N$-qubit Dicke battery: $\tau^{*}(\varepsilon)\geq\sqrt{N\varepsilon}/(2\lambda\sqrt{\bar{n}})$, valid for all $\lambda$, $N$, and $\varepsilon\in(0,1]$, and verified numerically for all $\bar{n}\geq 1$ with zero violations. The composite parameter $\Gamma_{N}=2\lambda\sqrt{\bar{n}/N}$ emerges naturally from the classical-field identity $(\hat{a}+\hat{a}^{\dagger})|\alpha\rangle=2\sqrt{\bar{n}}|\alpha\rangle$, which yields the collective spin-rotation frequency $\Omega_{N}=4\lambda\sqrt{\bar{n}}/\sqrt{N}$ and the exact short-time law $\varepsilon(t)=[1-\cos(\Omega_{N}t)]/2\approx(4/N)\lambda^{2}\bar{n}\,t^{2}$. The $\sqrt{\bar{n}}$ scaling reflects coherent enhancement: ergotropy charges $\bar{n}$ times faster than for a thermal charger of equal mean energy, in direct analogy with the Mandelstam–Tamm bound [22] applied to a thermodynamic observable. The QSL (5) translates into the circuit-QED design rule $\lambda\geq\sqrt{N\varepsilon}/(2\tau_{\rm target}\sqrt{\bar{n}})$ [20, 21]: for $\varepsilon=0.8$, $\tau_{\rm target}=100\,$ns, $\bar{n}=10$, $N=2$, one requires $\lambda/\omega_{0}\approx 0.020$, well within current experimental reach.
Three results distinguish this work. First, the exact short-time coefficient $A=4/N$ is derived analytically for arbitrary $N$, revealing that per-qubit ergotropy growth scales as $1/N$, a direct consequence of the $1/\sqrt{N}$ collective coupling in the Dicke Hamiltonian. Second, the global ergotropy bound $\varepsilon(t)\leq(4/N)\lambda^{2}\bar{n}\,t^{2}$ is proved analytically via $1-\cos x\leq x^{2}/2$ (Theorem 2), yielding a closed-form QSL for the Dicke battery. Third, the bound is tight at small $\varepsilon$: $\tau^{*}/\tau_{\rm QSL}\to 1^{+}$ as $\varepsilon\to 0$, confirmed over 7797 simulations with zero violations (Table 2). Unlike general observable speed limits [26], which bound rates of change without a closed-form expression for the Dicke model, our bound (5) is model-specific, analytically explicit in $(\lambda,\bar{n},N)$, and tight at small $\varepsilon$. While power bounds [27, 26] characterise the instantaneous charging rate, converting them into a minimum time for a prescribed ergotropy requires integrating a trajectory-dependent inequality — a step not addressed by existing frameworks. The commutation-matrix approach of Ref. [29] isolates the entanglement contribution to charging power in a model-independent way; our result complements this by providing an analytically explicit, tight bound for the Dicke model in terms of the experimentally accessible parameters $(\lambda,\bar{n},N)$. Natural extensions include Lindblad open-system dynamics [25, 18], analogous QSLs for the Tavis–Cummings and Jaynes–Cummings models, and experimental verification in circuit-QED.
Acknowledgments
A.J. gratefully acknowledges the support of the Department of Optics, Palacký University Olomouc, Czech Republic, where part of this work was carried out. A.E.A. completed part of this work during a research visit to the LPMMC, CNRS, in Grenoble, France. He extends his sincere gratitude to the CNRS – Fédération de Recherche QuantAlps, Comité de Direction QuantAlps, for their financial support and for fostering a stimulating and friendly research environment.